# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### 28 Feb 2026 v15

@@ 3.2. Quadratic Gradient: Definition
For the approximation $\tilde{B}$ to serve as a valid lower bound within the fixed-Hessian framework, it must satisfy the convergence criterion $\tilde{B}\leq H$. Given our foundational assumption that $\bar{H}\leq H$, this requirement is satisfied by establishing that $\tilde{B}\leq\bar{H}$. Following the analytical approach detailed in (IDASH2018bonte), it can be shown that $\tilde{B}$ consistently maintains this lower-bound property relative to $\bar{H}$, thereby ensuring stable convergence in the encrypted domain without the computational overhead of the full Hessian.
@@ 3.2.2. Systematic Fixed-Hessian Identification
Theoretically, the vanilla first-order gradient update, $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\alpha\boldsymbol{g}$ (where $\alpha$ is a sufficiently small step size), can be interpreted as a degenerate case of the fixed-Hessian Newton’s method. In this view, the inverse of the surrogate Hessian is simplified to a uniform scalar $\alpha\mathbf{I}$. However, such a scalar approximation fails to capture the non-uniform curvature of the objective function. Consequently, while it offers baseline stability, it lacks the adaptive convergence required for practical efficiency—particularly in high-dimensional encrypted optimization, where the high computational cost of each iteration necessitates a minimal number of steps to reach convergence.
@@ 3.3.1. Experiments
Empirically, for the standalone vanilla quadratic-gradient method, optimal performance is observed when the learning rate is situated within the interval $[1,2]$. Beyond a value of $3$, the algorithm typically exhibits divergent behavior. For the Enhanced NAG variant, we recommend an exponentially decaying learning rate to synchronize rapid early-stage progress with late-stage numerical stability.

### 16 Mar 2026 v16

@@ 3.2. Quadratic Gradient: Definition
For the approximation $\tilde{B}$ to serve as a valid lower bound within the fixed-Hessian framework, it must satisfy the convergence criterion $\tilde{B}\preceq H$. Given our foundational assumption that $\bar{H}\preceq H$, this requirement is satisfied by establishing that $\tilde{B}\preceq\bar{H}$. Following the analytical approach detailed in (IDASH2018bonte), it can be shown that $\tilde{B}$ consistently maintains this lower-bound property relative to $\bar{H}$, thereby ensuring stable convergence in the encrypted domain without the computational overhead of the full Hessian.
@@ 3.2.2. Systematic Fixed-Hessian Identification
From a theoretical standpoint, the standard first-order gradient update, $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\alpha\boldsymbol{g}$ (where $\alpha$ denotes a sufficiently small step size), can be interpreted as a degenerate instantiation of the fixed-Hessian Newton’s method. This equivalence holds under the condition that, with high probability, the scalar $\alpha$ is stochastically bounded by the diagonal elements of the Hessian approximation, i.e., $\alpha\leq\bar{B}_{kk}$ for all $k\in\{1,\dots,d\}$, where $\bar{B}_{kk}$ represents the $k$-th diagonal entry defined in the preceding Step 2. Consequently, the gradient descent update effectively serves as an isotropic approximation of second-order curvature, where the learning rate $\alpha$ implicitly regularizes the spectral properties of the approximated Hessian $\bar{\mathbf{B}}$. From a theoretical perspective, the vanilla first-order gradient update rule, expressed as $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\alpha\boldsymbol{g}$ (where $\alpha$ denotes a sufficiently small learning rate), can be interpreted as a degenerate case of the fixed-Hessian Newton’s method. This interpretation holds under the condition that, with high probability, the scalar $\alpha$ satisfies $\alpha\leq\bar{B}_{kk}$ for all diagonal elements $\bar{B}_{kk}$ of the Hessian approximation defined in the above step 2. Consequently, the standard gradient descent update can be viewed as an isotropic approximation of a second-order method, where the curvature information is bounded by the reciprocal of the step size. Theoretically, the vanilla first-order gradient update, $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\alpha\boldsymbol{g}$ (where $\alpha$ is a sufficiently small step size), can be interpreted as a degenerate case of the fixed-Hessian Newton’s method. In this view, the inverse of the surrogate Hessian is simplified to a uniform scalar $\alpha\mathbf{I}$. However, such a scalar approximation fails to capture the non-uniform curvature of the objective function. Consequently, while it offers baseline stability, it lacks the adaptive convergence required for practical efficiency—particularly in high-dimensional encrypted optimization, where the high computational cost of each iteration necessitates a minimal number of steps to reach convergence.
@@ 3.3.1. Experiments
Empirically, for the standalone vanilla quadratic-gradient method, optimal performance is observed when the learning rate is situated within the interval $[1,2]$. Beyond a value of $3$, the algorithm typically exhibits divergent behavior. For the Enhanced NAG variant, we recommend an exponentially decaying learning rate to synchronize rapid early-stage progress with late-stage numerical stability. Robustness and Generalization. Our proposed algorithm maintains consistent numerical stability and performance invariance even when the input space includes negative values, such as in datasets normalized to the interval [1,+1][-1,+1]. This sign-agnostic property ensures that the framework’s efficacy is independent of specific data scaling techniques. Furthermore, preliminary investigations suggest that our methodology can be extended beyond machine learning to broader numerical optimization problems. Consequently, our approach effectively addresses and mitigates two critical limitations inherent in the SFH framework: the constraint on non-negative input domains and the limited scope of application. By overcoming these bottlenecks, we significantly enhance the versatility and utility of the underlying optimization primitive.